You won the lottery and plan to buy your parents the coastal home they've always wanted. You pick out a $500,000 villa on the water, and make a $60,000 down-payment in cash. You take out a 30 year mortgage of $440,000 at 7.7% per year, and begin making the monthly mortgage payments.

After two years, your parents decide the house is too big and want to move.  

a) What is the unpaid balance on the mortgage after making two years of payments? (In other words, how much do you still owe on the house?)

b) How much total interest did you pay over two years?
Answer #1

a. Loan amount = 440000

i = 7.7% = 7.7% / 12 = 0.64167% per month

t = 30 * 12 = 360 months

Monthly loan payment = 440000 * (A/P, 0.64167%, 360)

= 440000 * 0.0064167*((1 + 0.0064167)^360)/((1 + 0.0064167)^360-1)

= 440000 * 0.0064167*((1.0064167)^360)/((1.0064167)^360-1)

= 440000 * 0.007130

= 3137.04

loan payments made after 2 yrs = 2* 12 = 24

Loan payments left = 360 - 24 = 336

Principal unpaid = 3137.04 * (P/A, 0.64167%, 336)

= 3137.04 * ((1 + 0.0064167)^336-1)/(0.0064167 * (1 + 0.0064167)^336)

= 3137.04 * ((1.0064167)^336-1)/(0.0064167 * (1.0064167)^336)

= 3137.04 * 137.674145

= 431888.87

b. Interest paid in two yrs = 3137.04 * 24 - (440000 - 431888.87) = 67177.76